A baseball diamond is 30yds each side. if u through a ball from second base to home plate how far would you through the ball.
Answer by nerdybill(7384)

A baseball diamond is 30yds each side. if u through a ball from second base to home plate how far would you through the ball.
.
Apply the Pythagorean Theorem:
Let x = distance between home plate and 2nd base
then
x^2 = 30^2 + 30^2
x^2 = 900 + 900
x^2 = 1800
x = sqrt(1800)
x = sqrt(10*10*3*3*2)
x = 30sqrt(2) yards (exact answer)
.
x = 42.43 yards
or
x = 127.28 feet